OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

i really need help on this one anyone please

OpenStudy (gorv):

n=3 f(3)=f(3-1)+f(3-2)=f(2)+f(1)=3=2=5

OpenStudy (gorv):

like this put values of n and calculate

OpenStudy (anonymous):

thanks

OpenStudy (amoodarya):

2, 3, 5, 8, 13, 21, 34, 55, 89, 144, 233, 377, 610, 987, 1597, 2584, 4181